OpenStudy (anonymous):

f^–1(x) teach me how to solve this

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

solve what? f^–1(x) equal: x/f ? fine

OpenStudy (anonymous):

domain= R (all real numbers)

OpenStudy (anonymous):

example: \[\LARGE f^{-1}(x)=\frac{2x-3}{x+1}\]

OpenStudy (anonymous):

inverse right?

OpenStudy (anonymous):

root: x = 0 f different 0

OpenStudy (anonymous):

inverse yes, i need to learn to solve inverse problems by tomorrow

OpenStudy (anonymous):

you have to replace all x-es with y and y with x... so you need to change their places and you'll have... from: \[\LARGE f^{-1}(x)=\frac{2x-3}{x+1}\] which is : \[\LARGE y=\frac{2x-3}{x+1}\] now change their places: \[\LARGE x=\frac{2y-3}{y+1}\] and now solve for y .... \[\LARGE x=\frac{2y-3}{y+1}\] \[\LARGE x(y+1)=2y-3\] \[\LARGE xy+x=2y-3\] \[\LARGE xy+x-2y=-3\] \[\LARGE xy-2y=-3-x\] \[\LARGE y(x-2)=-3-x\] \[\LARGE y=\frac{-3-x}{x-2}\]

OpenStudy (anonymous):

omg uhhh ok then but how do you know what f^-1(x)= in the first place